OpenStudy (anonymous):

Write 32in/6ft in simplest form

OpenStudy (anonymous):

Do you mean you want it in decimal form?

OpenStudy (anonymous):

@RyanL. no it asks for fraction

OpenStudy (anonymous):

I think that is as simple as it can get.

OpenStudy (anonymous):

nope its not it can go simpler cuz thats not even one of the answers i can choose from @RyanL.

ganeshie8 (ganeshie8):

hint- 1 foot = 12 inch

OpenStudy (ganpat):

1 ft = 12 inches.. so 32 in / 6 ft = 32 (6 * 12) = 8 / (6 * 3) = 4 / (3 * 3) = 4/9...

OpenStudy (ganpat):

is that in options ??

OpenStudy (anonymous):

thanks soo much!! i understand ow and yes it is @Ganpat

OpenStudy (ganpat):

w/c.. :)

OpenStudy (anonymous):

yes it is @Ganpat